TrigMC software release coordination: PTD's logbook - page for August 2018


Use reverse chronological order.

For each shift date show the NEW MRs, and any related relevant comments/questions; note down when I (or some one else) merged/closed each request, and mark the MR as done/closed by ticking the box ([x]). If an MR takes longer than a day to merge, leave it listed under the original shift date, but add comments prefaced with the date of the comment/action. Only MRs that are still active should have the box unticked ([ ]).

Therefore each MR will only appear once in the log, under the date of the day when I first became aware of it.

Requests still open:


[x] !13896 manual sweep from 12026 12027 12354 12285 (opened by Xiaohu Sun)

5.30pm This request was opened about 30 mins ago. However, it does not seem to be a proper merge request, as there are no actual changes required...? See this screenshot.

5/9/2018 10.44am This malformed MR has now disappeared. Good.

[x] !13879 Sweeping 13760 from 21.0 to 21.0-TrigMC. Fixing type problems in FTKWrapper outputs (opened by Atlas Nightlybuild)

10.57am MR was opened at 2am as usual. CI pipeline passed. Will approve, given it is a sweep. Merged OK.


[x] !13782 Update 21.0-TrigMC to 21.1.39 (opened by Rafal Bielski)

11.03am MR opened about 45 mins. ago. CI pipeline running.

4.50pm CI pipeline passed. Will wait for "review-approved" tag.

29/8/2018 12.23pm "Review-approved" label added 3h ago. Merged OK.

[x] Add EventOverlayJobTransforms-00-06-32-14 to (requested by Andrzej Olszewski)

10.37am Andrzej Olszewski emailed me a reminder that he needs the tag EventOverlayJobsTransforms-0-06-32-14 added to the current cache, and a release built out of it. Somehow I overlooked this (had email trouble while in Portugal last week...); original email was sent on Friday, just before I travelled + long weekend... Tag has now been added to the cache ( and is set to validate tonight.

29/8/2018 12.29pm Andrzej emailed confirming that he is happy with rel_3 (last night's nightly) after having run some tests. I have made some checks as well: cache compiled OK and all tests ran, with similar success fraction to the nightlies in this past week. Will ask for a build now.

29/8/2018 12.37pm Requested release out of the rel_3 cache, to the usual email address atlas-sw-release-shifters.

29/8/2018 4.21pm Oana Boeriu has made the new release and has confirmed that it is now finished; only awaiting Grid distribution.

29/8/2018 4.39pm Have emailed atlas-trig-relcoord and Andrzej to let them know that the new release is available, and that Grid distribution is in progress.


[x] !13685 Sweeping 13478 from 21.0 to 21.0-TrigMC. update FTKSim to match firmware in AUX track fitting (opened by Atlas Nightlybuild)

11.12am MR opened overnight. CI pipeline passed. Merged OK.


[x] !13633 Sweeping 13482 from 21.1-dev to 21.0-TrigMC. Adding back to the Trigger EDM the original ion jet trigger collection withoutů (opened by Atlas Nightlybuild)

8.47am Request opened at 2am. CI pipeline passed. Merged OK.


[x] !13609 Sweeping 13485 from 21.0 to 21.0-TrigMC. ATLASRECTS-4540: Add protection for invalidated vector iterators in SiTrackMaker_xk (opened by Atlas Nightlybuild)

11.27am Opened by Atlas Nightlybuild at the standard 2am slot (automatic sweep request)... CI pipeline concluded fine. Merged OK.


[x] !13531 Merge branch 'test_ART_TrigAnalysisTest_2' into '21.1-dev'. CP to 21.0-TrigMC (opened by Julie Hart)

4.11pm (FNAC) request opened 1h ago; pipeline has been running for 56m.

17/8/2018 10.44am Pipeline passed. "Review-approved" label added by Peter Berta about 1h ago. Merged OK.


[x] !13504 New MET trigger algorithm for FTK (21.0-TrigMC): track and topoclusters (ATR-16263) (opened by Renjie Wang)

6.08pm Request opened 10 mins ago. Pipeline is running.

16/8/2018 10.03am Pipeline passed 11h ago. Merged OK.


[x] !13466 Update 21.0-TrigMC to 21.0.78 (opened by Rafal Bielski)

9.08pm Pipeline has been running for 2h20m.

15/8/2018 11.59am Pipeline failed due to "required tests", about 12h ago. Rafal has made a commit about 2h ago to resolve the issue ("Resolving conflict between MC and Physics menu"). Pipeline is now running again, for 2h50 so far.

15/8/2018 3.24pm Pipeline passed a little while ago. As this is not a straight-forward sweep, and a new requester, will wait for review-approved label before proceeding with merge.

16/8/2018 9.59am "Review-approved" label has been added by Peter Berta. Merged OK.


[x] !13374 Sweeping 13348 from 21.1-dev to 21.0-TrigMC. Add HI_v5 tests to TriggerTest (opened by Atlas Nightlybuild)

12.15pm MR opened at 2am. CI pipeline OK and review-approved. Merged now: OK.


[x] !13212 Update 21.0-TrigMC to 21.1.37 (opened by Julie Hart Kirk)

1.15pm Request was opened a few hours ago. Pipeline started running 20 mins ago.

3/8/2018 10.35am CI pipeline passed yesterday evening. L1 approver has asked for L2 approval given extent of code changes for this MR. Will wait for that before merging.

4/8/2018 10.27am L2 approval was given yesterday afternoon. MR merged OK.

-- PedroTeixeiraDias - 02 Aug 2018

Edit | Attach | Watch | Print version | History: r22 < r21 < r20 < r19 < r18 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r22 - 20 Sep 2018 - PedroTeixeiraDias

This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2024 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback